OpenStudy (anonymous):

I can't figure out this one for some reason. (Probably missing a step or something. 0=361201-(p+1)^2

OpenStudy (anonymous):

First expand the (p+1)^2 like so (p+1)(p+1) After that it's arithmetic.

ganeshie8 (ganeshie8):

calculator is allowed or you have to figure out the root manually ?

OpenStudy (callisto):

\[0=361201-(p+1)^2\]\[361201-(p^2+2p+1)=0\]\[361201-p^2-2p-1=0\]\[-p^2-2p+361200=0\]\[p^2+2p-361200=0\]\[p=\frac{-2 \pm \sqrt{2^2 - 4(1)(-361200)}}{2(1)}\]\[p=\frac{-2 \pm \sqrt{1444804}}{2(1)}\]\[p=\frac{-2 \pm 1202}{2}\]Can you find the values of p?

OpenStudy (shubhamsrg):

why not subtract 361201 from both sides ,,cancel out the -ve sign..take sqrt of both sides..still gives you 2 ans..

OpenStudy (anonymous):

I have to find them manually. I thought that I was supposed to use the quadratic equation but it seemed somewhat difficult for me to do so I assumed there must be another way. Guess I'll have to remember to use it

OpenStudy (anonymous):

Thanks everyone!

OpenStudy (anonymous):

I meant formula, sorry. ^_^

ganeshie8 (ganeshie8):

(p+1) ^2 = 361201 \[p+1 = \pm \sqrt{361201}\] \[p + 1 = \pm 600... ?\] can you do trial and error few times to get the exact root..

OpenStudy (anonymous):

I think it would be best to leave the answer as a root. It's the most accurate.

ganeshie8 (ganeshie8):

agree.. but if it is a perfect square.. i would want to compute

OpenStudy (shubhamsrg):

sqrt would be 601,,and its not that difficult to know that..as you already know whats 600^2 601^2 = (600+1)^2 or 600^2 + 1201